OpenStudy (anonymous):

How do you integrate [(cosx)^4] (sinx)?

OpenStudy (anonymous):

let u=cosx du=??

OpenStudy (anonymous):

sinx

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

-sinx oops

OpenStudy (anonymous):

-sinxdx right then what?

OpenStudy (anonymous):

divide both sides of du=-sinxdx by -1 ?

OpenStudy (anonymous):

ok then?

OpenStudy (anonymous):

i replace sinxdx with -du

OpenStudy (anonymous):

hm i will solve it now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i just didn't know i had to use u subsittuion

OpenStudy (anonymous):

when do i use u substitution?

OpenStudy (anonymous):

replace cos^4x with u^4

OpenStudy (anonymous):

do i differentiate u^4 and (cosx)^4

OpenStudy (anonymous):

after substitution you have \[\int\limits_{}^{}-u ^{4}du\]

OpenStudy (anonymous):

should i use u=cosx or u^4 = (cosx)^4

OpenStudy (anonymous):

both are the same

OpenStudy (anonymous):

okay, i got what u wrote now

OpenStudy (anonymous):

i took out the negative 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then i integrated and replaced the u with cosx

OpenStudy (anonymous):

good

OpenStudy (anonymous):

omg yay i got the right answer! ty!

OpenStudy (anonymous):

have a good day

OpenStudy (anonymous):

thankyou :)